I need help on direct variation,finding constant of variation,then find value of y when x=3
Y=3 when x=-2
Thank you
I tried
Y=9\2. K=-3\2

Answer by ewatrrr(24785)

y = kx
3 = k2
3/2 = k
.....
y = (3/2)x
x = 3
y = 9/2